Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Friday, November 13, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 13, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/13/98 (Friday)

Segment 1 (Emidio to ANF). Report for 11/13/98. Excavation and Shoring of the pit at the culverts on Peace Valley Road , EMP 25.45, continued, the pit is over 24 feet deep. At the skip along Lebec Road, Potholing took place from ~EMP 21.75 to EMP 21.54. A dewatering structure, made of strawbales and a filter bag was constructed and water was pumped from the ROW for most of the day. At EMP 18.51, The welders cut-out a weld on the tie-in segment and tied-in again. NDT crews tested the tie-in welds. Excavation of tie-in trench continued at EMP 18.15. Shading and compaction took place on the Tejon Bench to EMP 17.28. Pipe is in the trench north to ~ EMP 17.02. Trench was excavated, with small skips from EMP 16.85 to EMP 16.60. Trenching continued northward on Grapevine road, to ~EMP 15.83. Pipe was welded at the top of Grapevine road to ~EMP 16.25. Work continued at Grapevine Station, with welding and electrical. Small crews worked at Emidio station. Backfill continued along Access Road B, where, the pipe is completely covered from Sabadan road to EMP 6.07. Backfill of the section from ~EMP .03 to ~EMP .55 continued. Excavation and Shoring continued in the area of EMP 1.

Segment 2 (ANF to Sylmar) Report for 11/13/98: 2A (Copco Rd to Templin Hwy): A drying 'pig' was pushed from the Copco Rd cutout, EMP 33.35, to 55.46 in Segment 2C. 2C (Templin Hwy to Castaic): Recontouring of the ROW at the Templin Hwy cutout, EMP 50.89 to 50.92, took place. All Project materials and equipment have been removed from the temporary storage area east of the I-5 bore, EMP 53.45. Erosion prevention materials were installed in gully crossings north of EMP 54.41. Revegetation and slope protection work has been completed on the steep slopes near EMP 54.74. Work began on stabilization of the spoils overrun at EMP 54.96. Shading reached north to the testing cutout at EMP 55.46. The contaminated soil at EMP 55.57 that resulted from yesterday's padding machine accident and resultant hydraulic fluid and motor oil spill has been collected for disposal. Recontouring continued near EMP 56.00. 2B (Castaic to Sylmar): At Feedmill Rd, the directional drill under Santa Clara River progressed 500 feet. The drill head was then retracted and replaced with a new bit. At Calgrove, the tie-in to the north side of the bore was completed with welding and x-ray. Also, more pipe was pre-fabricated at that location. Between Weldon Cyn. and Sierra Hwy, sawcutting in the roadway occurred, and 250 feet of trench was excavated. South of Sierra Hwy, bending and lowering-in of pipe moved north to Sierra Hwy. At Sunshine Cyn, 60 feet of F/O conduit was laid-in, 100 feet of slurry was applied, and an additional 100 feet of pipe was shaded with base rock through the intersection.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 13 November 1998: 3A: Crews continued working on the bore set up at Tuxford Street. Welding and x-ray worked on the north side tie in to the Hollywood Way Bridge crossing. Crews continued backfilling sink holes and assist pits along the MTA ROW near Tyburn and Fletcher Streets. 3B: No work occurred. 3C: Crews continued work for the aqueduct crossing.

Segment 4B(El Segundo to Lynwood) 11-13-98: The crew at El Segundo Station ESEMP 11.00 was working on the 90 degree tie-in. The crew at Western ESEMP 5.56 tied-in the last joint today. The paving crew was working at 108 Th St. ESEMP 1.16. The crew at Lynwood station was getting ready to send the cleaning pigs to the Corn field yard. The crew inside Lynwood station installed the power room and were working on the electrical wires and the back wall. The crew at College St. was repairing the fiberoptic conduit.

Segment 5 (Lynwood to Long Beach) Report for 13 November 1998: Reaming was completed at the directional drill at Highway 91/Compton Creek. Both cased bores at the Wilmington Refinery were completed and potholing for the bore pit on the east side of Alameda Street across from the station began.

Inspection Summary

In the past 24 hours, 5 inspection event(s) were reported for 5 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page